A few people do very poorly on a particular measure. What will be the effect on...

  1. A few people do very poorly on a particular measure. What will be the effect on the distribution of scores?

    a

    Positive skew

    b

    Negative skew

    c

    Bimodal

    d

    Platykurtic

  2. In a brief sentence describe how percentages are different than percentiles.

3. We have the following intelligence scores: 95, 100, 100, 100, 100, 105, 105.
What is the percentage of scores in the population that are above 100?

4.We have the following intelligence scores: 95, 100, 100, 100, 100, 105, 105.
If these data represented the population, what could we conclude about kurtosis?
Answer #1

1 Q. A few people do very poorly on a particular measure. What will be the effect on the distribution of scores?

A. egative skew Positive Sskew Bin eletykurticWhen one or a few scores are much higher or lower than all other scores, the curve is skewed.

Distribution is negatively skewed because few people scored much lower than all the other people

Answer is Negative Skew

2 Q. In a brief sentence describe how percentages are different than percentiles

A. A percentage is simply a representation of a proportion out of 100. To say three out of every six is the same as to say fifty out of every hundred, in other words fifty percent

A percentile is a statistical measure of distribution. For a given set of data, it is the level below which a certain percentage of the data falls. In graphic terms, it represents the area under the curve of a distribution.

For a more human example, if you score in the 72nd percentile on an exam, it means you scored higher than 72 percent of all the people who took the same exam— regardless of what your actual score was.

3 Q. We have the following intelligence scores: 95, 100, 100, 100, 100, 105, 105.
What is the percentage of scores in the population that are above 100?

A. Total number of Scores=6 (95,100,100,100,100,105,105)

Scores greater than 100=2 (105,105)

Percentage of scores in the population that are above 100=(scores greater than 100/total number of scores)*100

=(2/6)*100=0.33*100=33%

4 Q. We have the following intelligence scores: 95, 100, 100, 100, 100, 105, 105.
If these data represented the population, what could we conclude about kurtosis?

A. Kurtosis is one aspect of how scores are distributed

There are no extreme scores. This means that the curve is not skewed

Although not perfectly symmetrical, this distribution could be considered a normal, bell-shaped, mesokurtic curve